# Non-Gaussianities in primordial black hole formation and induced gravitational waves

### v2

@@ VI Conclusion and Discussion
Our starting point is the $\delta N$ formalism, which assumes that the Hubble patches we study are far enough from each other, such that each patch can be seen as a local Friedmann universe where the field value (and other perturbed quantities) takes a random initial value and evolves independently in the later stage. This is called seperate universe approach, which works quite well in the slow-roll regime. However, in the beginning of ultra-slow-roll stage (marked by $\varphi_{\mathrm{i}}$), the gradient terms of both growing mode and decaying mode are rather important, which give rise to the well-known $k^{4}$ growth Leach et al. (2001); Byrnes et al. (2019); Cole et al. (2022). Such a $k$-dependence cannot be recovered by seperate universe approach and the classical $\delta N$ formalism based on it Jackson et al. (2023). Fortunately, such a failure is only serious for $\varphi\lesssim\sqrt{\pi}\varphi_{\mathrm{i}}$ Domènech et al. (2023), which makes the seperate universe approach and the $\delta N$ formalsm still valid at the peak of the enhanced power spectrum, which is located at $\varphi_{\mathrm{p}}\approx\pi\varphi_{\mathrm{i}}$ Pi and Wang (2023). Apparently, if we extend the $\delta N$ formalism to include the spatial curvature on the initial slice taken slightly after horizon exit, the leading order correction of the gradient expansion can be taken into account Artigas et al. .
We did not discuss the interesting case of negative non-Gaussianity, which is connected to Type II PBHs. Simple estimate shows that negative non-Gaussianity will significantly suppress the PBH abundance, which can be calculated in a perturbative way Young and Byrnes (2013). For fully nonlinear curvature perturbation with a logarithmic dependence on the Gaussian variable, the PDF of $\mathcal{R}$ displays a Gumbel-like distribution Pi and Sasaki (2023b) (also called double exponential suppression Hooshangi et al. (2023)):
@@ Acknowledgement
I would like to thank Diego Cruces, Guillem Domènech, Jaume Garriga, Cristiano Germani, Cristian Joana, Misao Sasaki, Teruaki Suyama, Takahiro Tanaka, Dong-Gang Wang, and Jianing Wang for their comments and discussions, and to thank Ryoto Inui and Ao Wang for helping me draw Fig 7. I also thank the hospitality of YITP, Kavli IPMU, IBS-CTPU, and KIAS during my visit when this paper was finalized, and thank all the participants of the 2nd International Workshop on Gravitational Waves and the Early Universe for the intriguing discussions. This work is supported in part by the National Key Research and Development Program of China Grant No. 2021YFC2203004, by Project No. 12047503 of the National Natural Science Foundation of China, by the JSPS Grant-in-Aid for Early-Career Scientists No. JP20K14461, by JSPS KAKENHI No. JP24K00624, and by the World Premier International Research Center Initiative (WPI Initiative), MEXT, Japan.

### v3

@@ VI Conclusion and Discussion
Our starting point is the $\delta N$ formalism, which assumes that the Hubble patches we study are far enough from each other, such that each patch can be seen as a local Friedmann universe where the field value (and other perturbed quantities) takes a random initial value and evolves independently in the later stage. This is called seperate universe approach, which works quite well in the slow-roll regime. However, in the beginning of ultra-slow-roll stage (marked by $\varphi_{\mathrm{i}}$), the gradient terms of both growing mode and decaying mode are rather important, which give rise to the well-known $k^{4}$ growth Leach et al. (2001); Byrnes et al. (2019); Cole et al. (2022). Such a $k$-dependence cannot be recovered by seperate universe approach and the classical $\delta N$ formalism based on it Jackson et al. (2023). Fortunately, such a failure is only serious for $\varphi\lesssim\sqrt{\pi}\varphi_{\mathrm{i}}$ Domènech et al. (2023), which makes the seperate universe approach and the $\delta N$ formalsm still valid at the peak of the enhanced power spectrum, which is located at $\varphi_{\mathrm{p}}\approx\pi\varphi_{\mathrm{i}}$ Pi and Wang (2023). Apparently, if we extend the $\delta N$ formalism to include the spatial curvature on the initial slice taken slightly after horizon exit, the leading order correction of the gradient expansion can be taken into account Artigas et al. (2025).
We did not discuss the interesting case of negative non-Gaussianity, which is connected to Type II PBHs. Simple estimate shows that negative non-Gaussianity will significantly suppress the PBH abundance, which was calculated in a perturbative way Young and Byrnes (2013). For fully nonlinear curvature perturbation with a logarithmic dependence on the Gaussian variable, the PDF of $\mathcal{R}$ displays a Gumbel-like distribution Pi and Sasaki (2023b) (also called double exponential suppression Hooshangi et al. (2023)):
@@ Acknowledgement
I would like to thank Diego Cruces, Guillem Domènech, Jaume Garriga, Cristiano Germani, Cristian Joana, Misao Sasaki, Teruaki Suyama, Takahiro Tanaka, Dong-Gang Wang, and Jianing Wang for their comments and discussions, and to thank Ryoto Inui and Ao Wang for helping me draw Fig 7. I also thank the hospitality of YITP, Kavli IPMU, IBS-CTPU, and KIAS during my visit when this paper was finalized, and thank all the participants of the 2nd International Workshop on Gravitational Waves and the Early Universe for the intriguing discussions. This work is supported in part by the National Key Research and Development Program of China Grant No. 2021YFC2203004, by Project No. 12475066 and No. 12447101 of the National Natural Science Foundation of China, by the JSPS Grant-in-Aid for Early-Career Scientists No. JP20K14461, by JSPS KAKENHI No. JP24K00624, and by the World Premier International Research Center Initiative (WPI Initiative), MEXT, Japan.